A metal sphere of radius R is charged to a potential V. (a) Find the electrostatic energy stored in the electric field within a concentric sphere of radius 2 R. (b) Show that the electrostatic field energy stored outside the sphere of radius 2 R equals that stored within it.

Grade:10

1 Answers

Navjyot Kalra
9 years ago
Sol. Q = CV = 4πε base 0R * V W = 1/2 q^2/C [∴ ‘C’ in a spherical shell = 4 πε base 0R] E = 1/2 16π^2ε base 0^2 *R^2 * V^2/4πε base 0 * 2R = 2πε base 0RV^2 [‘C’ of bigger shell = 4 πε base 0R]